OpenStudy (anonymous):

b – 7 = 21 How do I work this out?

OpenStudy (anonymous):

add 7 to both sides so b=28

OpenStudy (anonymous):

b+7= 7 21 +7 = 28 Taht right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

b - 7 = 21 say that equation out loud : the number b minus seven is equal to twenty one what can b be? Can you think of a number such that if you subtract seven from it, you get 21? In other words, if you add 7 to 21 , you get such a number. that number is 28. you can verify it. twenty eight minus 7 is twenty one. therefore b is 28